OpenStudy (shamil98):

lim x-->0 f(x) = (e^x - 1)/(sin x) = 0/0 L'hopital's rule

OpenStudy (shamil98):

\[f(x) =\frac{ (e^x-1)' }{ (sinx)' }\]

OpenStudy (anonymous):

l'hospital rule states that if you have a lim that's 0/0 you can use f'x/g'x to get the limit

OpenStudy (anonymous):

l'hospital rule states that if you have a lim that's 0/0 you can use f'x/g'x to get the limit

OpenStudy (shamil98):

\[\frac{ e^x }{ \cos~x } = \frac{ e^0 }{ \cos 0 } = 1\] right? @Luigi0210

myininaya (myininaya):

1/1=1 is right

OpenStudy (anonymous):

That's right

OpenStudy (shamil98):

thanks

OpenStudy (anonymous):

welcome!